# with Self-Reflective Rationales

### v1

@@ 4.2 Main Experimental Results
Confidence Calibration Performance. We show the ECE results (Table 1) and the AUROC results (Table 2) to measure the correlation between the expressed confidence and the actual performance. We observe that SaySelf significantly outperforms all baseline approaches in reducing the calibration error (ECE) and improving the distinction of confidence in correct and incorrect responses (AUROC). This conclusion holds in both in-distribution (HotpotQA) and out-of-distribution datasets, which demonstrates the general applicability of SaySelf.
Task Performance. We show the accuracy results in Table 3. SC, which uses multiple sampling, achieves overall better performance compared to other approaches. However, this results in high inference latency. Compared to other baseline approaches, SaySelf can overall maintain the original task performance. This indicates that the task of confidence estimates doesn’t conflict with the original task, consistent with previous work [7, 47].
Faithfulness of the Generated Self-Reflective Rationales. The evaluation prompt for GPT-4 is shown in Appendix A. We show the faithfulness results in Table 4. Due to the budget limits for GPT-4 evaluation, we sample 100 instances from each dataset for evaluation. The instances with multiple selective reasoning chains are chosen for priority. For all baseline approaches, we explicitly instruct LLMs to provide reasoning for the confidence levels assigned to their prior responses. The results for DP and SC are combined, as both methods sampling from the same LLM.
@@ 4.3 Ablation Study
We conduct ablation study to verify several design choices in SaySelf: (1) w/o RL: We evaluate SaySelf without the reinforcement learning from the task supervision stage. (2) w/o R & CE: We evaluate SaySelf that directly train LLMs on the golden answer without the self-reflective rationales and confidence estimates in the supervised fine-tuning stage. (3) w/o R: We evaluate SaySelf that directly train LLMs on the golden answer and confidence estimates without the self-reflective rationales in the supervised fine-tuning stage. (4) Reward function: We verify the effectiveness of the defined reward function in SaySelf. We compare with a simple intuitive reward function: $R=I(\text{correct})\times\text{confidence level}-I(\text{incorrect})\times\text{confidence level}$.
The results are shown in Table 1, Table 2, and Table 3 for direct comparison with SaySelf. Compared with SaySelf w/o RL, our results indicate that while supervised fine-tuning can enable LLMs to express calibrated confidence to a certain extent, incorporating RL with task-specific supervision further enhances the accuracy of these confidence estimates. The ablation of the reward function also justifies our design choice in the RL stage. For the supervised fine-tuning stage, we find that both the self-reflective rationales and the confidence estimates contribute significantly to the calibrated confidence estimates. Overall, the ablation results verify the effectiveness of all components in SaySelf.
Figure 3: Case studies of SaySelf’s capability to generate insightful self-reflective rationales that effectively capture the internal uncertainty in LLMs. Various clusters illustrate a selection from 100 sampled responses, and the rationale is generated by LLMs.
@@ 4.4 Case Study
We perform a case study to better understand our approach (see Figure 3). We select 2 questions, prompt LLMs trained via SaySelf to generate the self-reflective rationales. Then we perform multiple sampling (100) and clustering to get a selection of representative responses. These two examples demonstrate that SaySelf exhibits strong capability in detecting and summarizing uncertainties internally. For example, in the first case, SaySelf expresses uncertainty about the exact location of the Howard Centre, identifying strong indications that it is likely in Letchworth and not Welwyn Garden City, with Cambridge being an unlikely option. This rationale acknowledges the mixed information leading to different founding years based on the location—1903 for Letchworth and 1920 for Welwyn Garden City, dismissing the 1841 Cambridge claim as highly improbable. This capability for self-reflective generation has profound impact on improving the reliability of LLM-based systems.

### v3

@@ 4.2 Main Experimental Results
Confidence Calibration Performance. We show the ECE results (Table 1) and the AUROC results (Table 5 in the Appendix) that measure the correlation between the expressed confidence and the actual performance. We observe that SaySelf significantly outperforms all baseline approaches in reducing the calibration error (ECE) and improving the distinction of confidence in correct and incorrect responses (AUROC). This conclusion holds in both in-distribution (HotpotQA) and out-of-distribution datasets, which demonstrates the general applicability of SaySelf. Also, the difference of SaySelf from other baselines is mostly statistically significant ($p<0.05$), further demonstrating its capability to provide effective confidence estimates.
Task Performance. We show the accuracy results in Table 2. SC, which uses multiple sampling, achieves overall better performance compared to other approaches. However, this results in high inference latency. Compared to other baseline approaches, SaySelf can overall maintain the original task performance. This indicates that the task of confidence estimates doesn’t conflict with the original task, consistent with previous work Chen et al. (2023b); Zhang et al. (2023a).
Faithfulness of the Generated Self-Reflective Rationales. The evaluation prompt for GPT-4 is shown in Appendix A. We show the faithfulness results in Table 3. Due to the budget limits for GPT-4 evaluation, we sample 100 instances from each dataset for evaluation. The instances with multiple selective reasoning chains are chosen for priority. For all baseline approaches, we explicitly instruct LLMs to provide reasoning for the confidence levels assigned to their prior responses. The results for DP and SC are combined, as both methods sample from the same LLM.
@@
We conduct human annotations to justify the use of GPT-4 for automatic evaluation. The details are described in Appendix C. We observe a Spearman’s rank correlation coefficient of 0.89 between the ratings given by GPT-4 and humans, which demonstrates the reliability of automatic evaluation.
@@ 4.3 Ablation Study
We conduct an ablation study to verify several design choices in SaySelf: (1) w/o RL: We evaluate SaySelf without the reinforcement learning from the task supervision stage. (2) w/o R & CE: We evaluate SaySelf that directly trains LLMs on the golden answer without the self-reflective rationales and confidence estimates in the supervised fine-tuning stage. (3) w/o R: We evaluate SaySelf that directly trains LLMs on the golden answer and confidence estimates without the self-reflective rationales in the supervised fine-tuning stage. (4) w/ Naive RF: We verify the effectiveness of the defined reward function in SaySelf. We compare with a simple intuitive reward function: $R=\mathbb{I}(\text{correct})\times\text{confidence level}-\mathbb{I}(\text{incorrect})\times\text{confidence level}$.
The results are shown in Table 1, Table 6, Table 5 (Appendix), and Table 2 for direct comparison with SaySelf. Compared with SaySelf w/o RL, our results indicate that while supervised fine-tuning can enable LLMs to express calibrated confidence to a certain extent, incorporating RL with task-specific supervision further enhances the accuracy of these confidence estimates. The ablation of the reward function also justifies our design choice in the RL stage. For the supervised fine-tuning stage, both the self-reflective rationales and the confidence estimates contribute significantly to the calibrated confidence estimates. Overall, the ablation results verify the effectiveness of all components in SaySelf.
Figure 3: Case study of SaySelf’s capability to generate insightful self-reflective rationales that effectively capture the internal uncertainty in LLMs. Various clusters illustrate a selection from 100 sampled responses, and the rationale is generated by LLMs. Another example is given in Figure 4 in the Appendix.
Unanswerable Answerable $\ \ \ \Delta$
DP 0.6696 0.7646 0.0950
SC 0.2317 0.3561 0.1244
PC 0.8676 0.9974 0.1298
R-Tuning 0.7614 0.8210 0.0596
AS 0.4381 0.4503 0.0122
GCE 0.2987 0.2991 0.0004
SaySelf 0.4962 0.7406 0.2444
Table 4: The confidence in unanswerable and answerable subsets of SQUADRUN.
4.4 Unanswerable Questions
We measure whether LLMs demonstrate low confidence in responding to unanswerable questions, which serves as a clear indicator of their ability to accurately delineate their knowledge boundaries We choose the SQUADRUN dataset (Rajpurkar et al., 2018), which contains both answerable and unanswerable questions. We measure the average confidence in the answerable and unanswerable subsets (see Table 4). We show that SaySelf enables LLMs to significantly reduce the confidence in unanswerable questions while maintain the confidence in the answerable parts, achieving the best confidence gap ($\Delta$) between the two subsets.
@@ 4.5 Case Study
We perform a case study to better understand our approach (see Figure 3 and Figure 4 in the Appendix). We choose two straightforward questions from HotpotQA and prompt LLMs trained via SaySelf to generate the self-reflective rationales. Then we perform multiple sampling (100 times) and clustering to get a selection of representative responses. These examples demonstrate SaySelf ’s strong ability to detect and summarize internal uncertainties. For example, in the first case, SaySelf expresses uncertainty about the exact location of the Howard Centre, identifying strong indications that it is likely in Letchworth and not Welwyn Garden City, with Cambridge being an unlikely option. This rationale acknowledges the mixed information leading to different founding years based on the location—1903 for Letchworth and 1920 for Welwyn Garden City, dismissing the 1841 Cambridge claim as highly improbable. This capability for self-reflective generation has a profound impact on improving the reliability of LLM-based systems.
@@ 5 Conclusion
Limitations
A potential limitation of SaySelf is its dependence on multiple sampled chains of reasoning to develop self-reflective rationales for training. There is still an ongoing debate regarding the faithfulness of CoT reasoning, specifically questioning whether it authentically represents the thinking process of LLMs (Lanham et al., 2023; Bentham et al., 2024; Turpin et al., 2024). The unfaithful CoT reasoning can cause unfaithful self-reflective rationales. Nonetheless, our ablation study demonstrates that these self-reflective rationales substantially enhance calibration performance. Further improvements in the effectiveness and faithfulness of SaySelf could potentially be achieved by integrating methods from recent research aimed at increasing the faithfulness of CoT reasoning, as suggested by Lyu et al. (2023).
Ethical Considerations
This work aims to improve the performance of LLMs in eliciting more fine-grained confidence estimates and self-reflective rationales. In the case of this work, it involves the use of Mistral 7B and GPT-4, so the same risks from LLMs research are also applicable to this work Bender et al. (2021).
While SaySelf aims to enhance trust in AI by providing clear confidence expressions and self-reflective rationales, there is a risk that users might over-rely on these confidence estimates. If the self-reflective rationales are not accurate or fail to capture the true uncertainty of the model, it could lead to potentially harmful decisions based on the model’s outputs. Therefore, users are advised to check important information before making crucial decisions.
This paper works on several publicly available datasets including HotpotQA, TruthfulQA, StrategyQA, FEVER, HaluEval, and ParaRel. They are available for the research community to study under Apache 2.0, Apache 2.0, MIT, CC-BY-SA 3.0, Apache 2.0, and MIT licenses respectively. Data is anonymized, thus our work does not propagate any privacy problems about any specific entities.
Finally, we carried out human annotations for analysis purposes. Since the amount of work is small, we and the annotators agree to consider it as a voluntary service. We have sufficiently discussed the specific use of the annotations and potential risks to annotators before the work.
Acknowledgments
This work is supported in part by the US National Science Foundation under grant NSF-IIS2226108. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 GPT-4 Summarization Prompt
DP 0.3222 0.5667 0.5193 0.5371 0.5278 0.5291
SC 0.5765 0.4939 0.5498 0.5472 0.5843 0.5546
PC 0.6636 0.5244 0.5385 0.5139 0.5544 0.6181
R-Tuning 0.6529 0.5980 0.5406 0.5688 0.5330 0.5424
AS 0.4955 0.4835 0.5391 0.5101 0.5569 0.5713
GCE 0.5042 0.4966 0.5043 0.4942 0.4907 0.5031
SaySelf 0.7156 0.6107 0.6074 0.6481 0.7318 0.6816
w/o RL 0.6524 0.5675 0.5910 0.5798 0.5929 0.6003
w/o R & CE 0.5256 0.5724 0.5738 0.6059 0.6002 0.5823
w/o R 0.4928 0.4952 0.4567 0.4831 0.4893 0.4853
w/ Naive RF 0.5140 0.4907 0.5091 0.5137 0.5147 0.5053
Table 5: The AUROC evaluation results of baselines, SaySelf, and various ablations.
Method | Dataset HotpotQA FEVER HaluEval
ECE Accuracy ECE Accuracy ECE Accuracy
DP 0.4744 0.4762 0.6167 0.6282 0.7750 0.8884
SC 0.4798 0.3475 0.4775 0.6325 0.5220 0.7168
PC 0.4604 0.3678 0.4513 0.6062 0.3375 0.8159
R-Tuning 0.4094 0.4348 0.4250 0.5588 0.3393 0.6604
AS 0.3525 0.4230 0.3944 0.4937 0.3027 0.9150
GCE 0.4056 0.3480 0.4330 0.5860 0.3190 0.8650
SaySelf 0.3296 0.4284 0.3844 0.6125 0.2427 0.8552
Table 6: The ECE and accuracy evaluation results of baselines and SaySelf on a different base model, Llama 3 8B.
Figure 4: Case study of SaySelf’s capability to generate insightful self-reflective rationales.
@@ A.2 GPT-4 Evaluation Prompt
Appendix B Empirical Evidence
In our study, we conduct both quantitative and qualitative analyses of the diversity of reasoning chains within each clustering group. The quantitative analysis reveals an average similarity of 0.94 across the reasoning chains, indicating high consistency and similarity within each cluster. For qualitative analysis, we randomly pick 3 clusters and manually inspect the reasoning chains in each cluster. We discover that the exact similarity rates within these clusters are 58%, 80%, and 74%, respectively, with the variations primarily involving minor differences in wording and sentence structure in the remaining reasoning chains. This supports our design decision to select one instance per cluster at random.
Appendix C Human Annotations for GPT-4 Evaluation
We randomly select 200 questions from multiple test datasets, providing each question along with corresponding reasoning chains and self-reflective rationales to two human annotators. The two human annotators are PhD students. These annotators, guided by the instructions for GPT-4, are asked to rate each case on a scale from 1 to 10, and the final score for each case is the average over two human annotators. We measure the pearman’s rank correlation coefficient between human evaluation and GPT-4 evaluation. The correlation coefficient is 0.89, which proves the reliability of automatic evaluation using GPT-4.
Appendix D Experiments of SaySelf on Llama
To test the generalizability of SaySelf, we test SaySelf on a different base model, Llama 3 8B. The baselines are the same as mentioned in Section 4.1. The experiment results are given in Table 6. Our results indicate that SaySelf can generalize in different base models and has superior performance over other baseline methods in ECE, without a significant loss in accuracy.